OpenStudy (tennisthegreat123):

What is the numerator of the fraction 4/15 that has a denominator of 60? A. 6 B. 8 C. 12 D. 16

OpenStudy (destinykiara99):

D. 16

OpenStudy (tennisthegreat123):

Can you tell me how to do it? I dont understand

OpenStudy (destinykiara99):

Yeah. Start off with 4/15 and set it up as : 4/15=x/60. 60/15=4. Multiply 4*4 and get 16 for the numerator.

OpenStudy (tennisthegreat123):

@destinykiara99 That doesnt make sense at all :P

OpenStudy (destinykiara99):

Do you know how to cross multiply?? That is the easiest way to learn how to do that type of problem

OpenStudy (tennisthegreat123):

no not really. i can ask my teacher tomorrow. i just have one more similar problem

OpenStudy (tennisthegreat123):

What is the numerator of the fraction equivalent to 24/120 that has a denominator of 20?

OpenStudy (tennisthegreat123):

A. 4 B. 6 C. 12 D. 16

OpenStudy (destinykiara99):

I can help. One sec I will try to write it out.

OpenStudy (destinykiara99):

A. 4

OpenStudy (tennisthegreat123):

ok

OpenStudy (tennisthegreat123):

thanks

OpenStudy (destinykiara99):